OpenStudy (anonymous):

write the equation of the line that passes through (-5,8) and (-3,-8)

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(-8-8)/(-3-(-5)) m=(-16)/(-3-(-5)) m=(-16)/(-3+5) m=(-16)/(2) m=-8 So the slope of the line that goes through the points is m=-8 Now let's use the point slope formula: y-y1=m(x-x[1]) y-8=-8(x--5) y-8=-8(x+5) y-8=-8x-8(5) y-8=-8x-40 y=-8x-40+8 y=-8x-32 So the equation of the line that goes through the two points is y = -8x-32